# DPER: Diffusion Prior Driven Neural Representation for Limited Angle and Sparse View CT Reconstruction

### v1

@@ 5 Conclusion
This paper proposes DPER, an unsupervised DL model for highly ill-posed LACT and ultra-SVCT tasks. The proposed DPER follows the HQS framework to decompose the classical CT inverse problem into a data fidelity sub-problem and a distribution prior sub-problem. Then, the two sub-problems are respectively addressed by INR and a pre-trained diffusion model. The INR guarantees data consistency with incomplete CT measurements benefiting from the integration of the CT physical forward model, while the diffusion model provides a generative image prior effectively constraining the solution space. The empirical studies on two datasets reveal that the proposed DPER excels both quantitatively and qualitatively, establishing itself as the state-of-the-art performer on in-domain and out-of-domain datasets for the LACT problem, while for the ultra-SVCT problem, it demonstrates performance that is on par with other SOTA methods.

### v2

@@ 5 Conclusion
This paper proposes DPER, an unsupervised DL model for highly ill-posed LACT and ultra-SVCT tasks. The proposed DPER follows the HQS framework to decompose the classical CT inverse problem into a data fidelity sub-problem and a distribution prior sub-problem. Then, the two sub-problems are respectively addressed by INR and a pre-trained diffusion model. The INR guarantees data consistency with incomplete CT measurements benefiting from the integration of the CT physical forward model, while the diffusion model provides a generative image prior effectively constraining the solution space. The empirical studies on comprehensive comparison experiments reveal that the proposed DPER significantly excels both quantitatively and qualitatively. It establishes itself as the state-of-the-art performer on in-domain and out-of-domain datasets for both the LACT and ultra-SVCT problems.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) under Grant 62071299 and Grant 12074258.